A record turntable rotating at 33.33 rev/min slows down and stops in 30 sec. after the motor is turned off.
How would i find its constant angular acceleration in rev/min^2
1
Expert's answer
2012-07-19T09:09:34-0400
Let a be the angular acceleration. Then

W = aT,

where T is deceleration time and W is initial angular velocity. Taking into account that 30 sec = 0.5 min, we'll get

a = W/T = 33.33/0.5 ≈ 66.66 rev/min².